# Deep Synoptic Array Science: Searching for Long Duration Radio Transients with the DSA-110

### v2

@@ I Introduction
Radio pulsar searching is entering a new exploratory phase; traditional single-dish telescope searches such as the Pulsar Arecibo L-band Feed Array (PALFA), Parkes Multi-beam Survey (PMBS), and Five-Hundred Metre Aperture Spherical Telescope (FAST) have utilized multi-beam receivers to discover roughly 2/3 of the $\sim 4000$ known Galactic pulsars (manchester2005australia; manchester2001parkes; han2021fast; li2025searching). Over the past decade, the TRAnsients and PUlsars with MeerKAT (TRAPUM; carli2024trapum), the Canadian Hydrogen Intensity Mapping Experiment Pulsar (CHIME/PULSAR; dong2023second) backend, the LOw-Frequency ARray (LOFAR) Tied-Array All-Sky Survey (LOTAAS van2023lofar), and Very Large Array (VLA) realfast system (law2018realfast), among others, have demonstrated that radio interferometers can achieve competitive sensitivities to single-dish telescopes with higher localization precision and less-costly antennas. Beyond contributing $\sim 500$ additional pulsars, these arrays have been essential for follow-up imaging to identify and associate Supernova Remnants (SNRs), Pulsar Wind Nebulae (PWNe), long-duration afterglows, and Persistent Radio Sources (PRS) of both pulsars and extragalactic radio transients (e.g. gaensler2005expanding; bruni2025discovery; cotton2024meerkat). However, these surveys have typically targeted pulsars with rotation periods $\lesssim 10$ seconds out of necessity. Fast Fourier-domain search strategies require long integration times to remain sensitive to periods beyond second-scales, and suffer from low-frequency red noise, making them impractical targets (ransom2001new; singh2022gmrt). Periodic transients with longer rotation periods (seconds to minutes), such as radio magnetars, Rotating Radio Transients (RRATs), and White Dwarf-M Dwarf binaries (either non-tidally-locked intermediate polars like AR Scorpii or tidally-locked non-accreting polars) were instead identified through serendipitous single-pulse surveys or follow-up of their high-energy counterparts (e.g. hyman2005powerful; caleb2022radio; marsh2016radio).
@@
The recent discovery of a dozen Galactic radio sources with unusually long pulse periods between $\sim 20$ seconds8.7-8.7 hours and low $\lesssim 10\%$ duty cycles have begun to reveal an abundant unconstrained parameter space (e.g. hurley2022radio; de2025sporadic; horvath2025unified; men2025highly; tan2018lofar). These so-called “Long-Period Radio Transients” (LPRTs, or LPTs) exhibit some properties characteristic of neutron stars, such as high polarization fractions, polarization position angle (PA) swings, temporal sub-structure, and un-detectable optical/IR (OIR) emission (mitra2023meterwavelength; kramer2024quasi; 1969ApL.....3..225R; chrimes2022new). However, at least two LPRTs, ILT J1101+5521 ($P=2.1$ hours) and GLEAM-X J0704-37 ($P=2.9$ hours) reside in binaries with M-dwarf companions, with blue optical excess suggesting they are White Dwarfs (de2025sporadic; hurley20242). While the remaining sample have no clear M-Dwarf companions, the recent discovery of a secondary $P=8.74$ hr period of GPM J1839-10 solely from radio timing suggests the White-Dwarf-M-Dwarf interpretation may extend to OIR-quiet candidates (horvath2025unified; men2025highly). Nonetheless, a plethora of both neutron star and White Dwarf-M Dwarf binary models have been proposed to explain these new phenomena (e.g. beniamini2023evidence; beniamini2020periodicity; cooper2024beyond; qu2025magnetic).
@@
Perhaps most promising is the discovery of four ASKAP LPRTs (ASKAP J1755-2527, ASKAP J1935+2148, ASKAP J1839-0756, ASKAP/DART-J1832-0911), LOFAR’s ILT J1101+5521, and the MWA’s GPM J1839-10 and GLEAM-X J0704-37 using second-duration radio image differencing pipelines (wang2025detection; li202444; lee2025emission; caleb2024emission; de2025sporadic; mcsweeney2025new; dobie2024two; hurley2023long). Radio imaging consists of binning the U-V (East-West and North-South coordinates of each antenna pair’s separation) visibility coordinates on a grid and applying a 2-dimensional FFT. This is much faster than visibility beamforming and is mathematically identical, though some resolution and flux sensitivity are lost via the gridding process. The image-plane LPRT detections demonstrate that this sensitivity loss is negligible for LPRT timescales (seconds-to-minutes), motivating the development of real-time LPRT image plane search pipelines like the Commensal Real-time ASKAP Fast-Transients Coherent (CRACO) Upgrade (wang2025craft), in addition to upcoming beamformer-based searches such as the CHIME/SLOW pipeline (J. Huang et al., in prep.).
In this paper, we introduce the Deep Synoptic Array’s (DSA-110) radio telescope’s Not-So-Fast Radio Burst (NSFRB) pipeline, an image-plane single-pulse search for LPRTs on 0.134-160.8 second timescales. The NSFRB system builds on the existing real-time Fast Radio Burst (FRB) search, which has discovered $\sim 60$ FRBs and localized $\sim 30$ to host galaxies with arcsecond-scale precision (ravi2023deep; law2024deep; sharma2024preferential). In Section II we describe the NSFRB system design and search pipeline; in Section III we describe the commissioning NSFRB injection tests and observations of pulsar B0329+54; in Section IV we report the results of a pilot DSA-110 NSFRB Galactic Plane Survey (DN-GPS); in Section V we discuss the sensitivity limits placed on LPRT emission in the Galactic Plane, and its implications for the White Dwarf binary model; in Section VI we conclude with next steps, specifically the real-time search which is currently underway.
@@ II NSFRB System Design
The DSA-110 is a 96-antenna drift-scan (see e.g. mantovanini2025galactic; 2015PASA...32...25W, for examples of drift-scan transient surveys with the MWA) radio interferometer at the Owens Valley Radio Observatory (OVRO) in Big Pine, CA. Details on the 64-antenna deployment and commissioning FRB survey can be found in ravi2023deep. The array was completed in early 2025 by supplementing the East-West row of 47 “core” antennas with a 35-antenna North-South arm, enabling real-time localization of radio sources without incorporating the 14 long-baseline “outrigger” antennas. The antenna design, RF signal chain, baseband processing, and general compute resource layout remain largely similar to the 64-antenna deployment, and will not be detailed here. The revised visibility correlator, bandpass calibration procedures, and the FRB search pipeline will be discussed in detail elsewhere (Ravi et al., in prep.) For this work, we focus on components specific to the NSFRB system. Code for the NSFRB search is public at https://github.com/dsa110/dsa110-nsfrb, all of which is implemented using Python, leveraging the JAX [^2] https://docs.jax.dev/en/latest/index.html library to deploy GPU tasks. An ETCD [^3] https://etcd.io/ server is used to manage alerts and coordination between compute nodes. Figure 1 shows the full real-time system diagram, which we describe in the next sub-sections.
@@ II.1 Fringe-Stopping and Imaging
for the DSA-110, $b_{\rm max}=2.6\,$km, $\lambda_{c}=20$ cm, where $\delta$ is the declination (e.g. thompson2017interferometry). The array operates in drift-scan mode (no azimuthal slew) at $\delta=71.6^{\circ}$ during normal operations, making the nominal fringe rate $t_{F0}\approx 3.36\,$s.
Fringe-stopped visibilities are written to a 14.9 MB buffer read by the correlator node imager, which grids visibilities with baseline lengths $b>b_{\rm min}=20\,$m from the dense 82-antenna ‘core’ ($b<b_{\rm max}=485\,$m) to a $175\times 175$-pixel UV-grid. The core is assumed to be coplanar, so the W-term is neglected for normal operation; however, W-stacking is implemented and may be explored in the future (e.g. gheller2023high). Outrigger antenna data is included only for detailed manual inspection of candidates. The two visibility polarizations are summed, then approximately uniform weighting (Briggs robustness parameter $\mathcal{R}=-2$; briggs1995high) is applied to accumulate visibilities, and a 2-dimensional inverse-FFT [^4] The NumPy Python Package implementation of np.fft.ifft2d is used for imaging. The visibility grid is first shifted using np.fft.ifftshift to properly order the zero-frequency component; np.fft.ifftshift is then applied to the resulting image. is applied to form the final image with $\Delta\alpha=10.3^{\prime\prime}(\cos{\delta})^{-1}\,,\,\Delta\delta=10.3^{\prime\prime}$ pixel scale (3 pixels per $31^{\prime\prime}$ synthesized beamwidth). The DSA-110 primary beam FWHM is roughly $\Omega=1.5^{\circ}(\cos{\delta})^{-1}\times 1.5^{\circ}$, while the $175\times 175$ pixel image covers a $\Omega=0.5^{\circ}(\cos{\delta})^{-1}\times 0.5^{\circ}$ search region. Each of the 8 channels are imaged individually, then summed together to form a single image for each 134134\,ms time sample. Images on each node are then buffered (6.125 MB) for transport by an http client to a central search server.
@@
The custom cerberus single-pulse search pipeline implements Python JAX Just-In-Time (JIT) compiled functions to de-disperse and boxcar filter images on the GPU. Data is first median subtracted in each pixel, then de-dispersed over 16 DM trials between 1714000171-4000 pc cm-3 (DM tolerance of 1.6). The JAX de-dispersion method is adapted from a custom PyTorch implementation, PyTorchDedispersion [^5] https://github.com/nkosogor/PyTorchDedispersion (2019arXiv191201703P; 2025ApJ...985..265K), and allows shift indices for each frequency channel to be pre-computed and loaded onto the GPU to minimize latencies. The resulting de-dispersed time series are boxcar filtered on $\log_{2}$ spaced trials from 1-16 samples (0.1342.140.134-2.14 s), then normalized by the running standard deviation to compute the S/N. Final candidates are found by taking the maximum S/N (and its TOA) over time. A nominal $6\sigma$ signal-to-noise (S/N) threshold is imposed to identify candidates, which are pushed to the post-processor.
The search is performed at three sampling timescales: 0.134 s, 0.67 s (binned by 5 samples), and 10.05 s (binned by 75 samples). The process server bins images at each timescale as they are received, aligning them in the RA direction to account for the 25-sample fringe-stopping interval, and preserves batches for cerberus to re-search at the 0.67 s and 10.05 s timescales. For the real-time search, a single search iteration on the next shortest timescale (0.134 s, 0.67 s) is dropped to enable the longer timescale (0.67 s, 10.05 s) search iterations (the Galactic Plane data is not searched in real-time and thus does not drop iterations). The 0.67 s data is searched identically to the 0.134 s data, now probing DMs from 85520000855-20000\,pc cm-3666While it is not strictly necessary to search up to 2000020000\,pc cm-3 since the peak DMs in the Galactic Plane typically only reach $\sim 5000$pc cm-3, this range was selected because (1) the coarse time and frequency resolution, there is little benefit from finely sampled DM trials, and (2) extending the DM range by exactly $5\times$ for data sampled at $5\times$ lower resolution allows the GPU search code to utilize the same pre-computed shifts, thus saving GPU compute time and memory by only loading the the lookup table once. with widths from 0.6710.720.67-10.72\,s. The longest timescale data bypasses de-dispersion, leveraging the fact that the DM delay for DMs below $\sim 20000\,$pc cm-3 is now contained to a single sample across the 187.5 MHz band. They are then searched from 10.05160.810.05-160.8 s. The sensitivity of the search is shown in Figure 2, which indicates the full range of single-pulse widths the NSFRB search probes. The broad morphology of known LPRTs includes those with smooth profiles (e.g. mcsweeney2025new) and ‘spiky’ sub-pulses (e.g. li202444; dong2025chime). By searching at multiple timescales, cerberus will be sensitive to both morphologies.
@@
Figure 2: Pulse width and $6\sigma_{\rm srch}$ Fluence Sensitivity to $\Delta t=134$ ms Bursts of the DSA-110 FRB (light blue; $\sim 840\,$mJy ms) and NSFRB (dark blue hatched; $\sim 40\,$Jy ms) searches. For each search we assume the fluence threshold is fixed; thus the lower border indicates the corresponding flux density threshold $6\sigma_{\rm th}\Delta t/W$ where $W$ is the width measured at 1.4 GHz (note that in some cases (e.g. ASKAP J1832-09 li202444; wang2025detection), only sub-pulses were detected at 1.4 GHz while a broader profile was detected at lower frequencies. In these cases, we still use the 1.4 GHz sub-pulse width). The dark blue dashed line shows the 90% completeness limit $25\sigma\approx 1200$ mJy ($25\sigma\Delta t\approx 160$ Jy ms; see Section III.2). The purple dashed line shows the theoretical radiometer $6\sigma_{\rm th}$ sensitivity of the NSFRB survey. 1.4 GHz fluences and widths are shown for radio pulsars in the ATNF catalog (diamonds), FRBs from the literature (triangles), and LPRTs (circles). DSA-110 FRBs have dark blue outlines and show their detection pulse widths and S/N rather than refined estimates from baseband analysis. The colorbar indicates distance estimates in kpc (FRB luminosity distances are computed from host Galaxy redshifts; pulsar, magnetar, and LPRT distances are from parallaxes where available, and DM distance limits for the rest of the sample). Note that only pulsars, FRBs, and LPRTs with 1.4 GHz detections are shown.

### v3

@@ I Introduction
Radio pulsar searching is entering a new exploratory phase; traditional single-dish telescope searches such as the Pulsar Arecibo L-band Feed Array (PALFA), Parkes Multi-beam Survey (PMBS), and Five-Hundred Metre Aperture Spherical Telescope (FAST) have utilized multi-beam receivers to discover roughly 2/3 of the $\sim 4000$ known Galactic pulsars (Manchester et al., 2005, 2001; Han et al., 2021; Li et al., 2025). Over the past decade, the TRAnsients and PUlsars with MeerKAT (TRAPUM; Carli et al., 2024), the Canadian Hydrogen Intensity Mapping Experiment Pulsar (CHIME/PULSAR; Dong et al., 2023) backend, the LOw-Frequency ARray (LOFAR) Tied-Array All-Sky Survey (LOTAAS van der Wateren et al., 2023), and Very Large Array (VLA) realfast system (Law et al., 2018), among others, have demonstrated that radio interferometers can achieve competitive sensitivities to single-dish telescopes with higher localization precision and less-costly antennas. Beyond contributing $\sim 500$ additional pulsars, these arrays have been essential for follow-up imaging to identify and associate Supernova Remnants (SNRs), Pulsar Wind Nebulae (PWNe), long-duration afterglows, and Persistent Radio Sources (PRS) of both pulsars and extragalactic radio transients (e.g. Gaensler et al., 2005; Bruni et al., 2025; Cotton et al., 2024). However, these surveys have typically targeted pulsars with rotation periods $\lesssim 10$ seconds out of necessity. Fast Fourier-domain search strategies require long integration times to remain sensitive to periods beyond second-scales, and suffer from low-frequency red noise, making them impractical targets (Ransom, 2001; Singh et al., 2022). Periodic transients with longer rotation periods (seconds to minutes), such as radio magnetars, Rotating Radio Transients (RRATs), and White Dwarf-M Dwarf binaries (either non-tidally-locked intermediate polars like AR Scorpii or tidally-locked non-accreting polars) were instead identified through serendipitous single-pulse surveys or follow-up of their high-energy counterparts (e.g. Hyman et al., 2005; Caleb et al., 2022a; Marsh et al., 2016).
@@
The recent discovery of a dozen Galactic radio sources with unusually long pulse periods between $\sim 20$ seconds8.7-8.7 hours and low $\lesssim 10\%$ duty cycles have begun to reveal an abundant unconstrained parameter space (e.g. Hurley-Walker et al., 2022; de Ruiter et al., 2025; Horváth et al., 2025b; Men et al., 2025; Tan et al., 2018). These so-called “Long-Period Radio Transients” (LPRTs, or LPTs) exhibit some properties characteristic of neutron stars, such as high polarization fractions, polarization position angle (PA) swings, temporal sub-structure, and un-detectable optical/IR (OIR) emission (Mitra et al., 2023; Kramer et al., 2024; Radhakrishnan and Cooke, 1969; Chrimes et al., 2022). However, at least two LPRTs, ILT J1101+5521 ($P=2.1$ hours) and GLEAM-X J0704-37 ($P=2.9$ hours) reside in binaries with M-dwarf companions, with blue optical excess suggesting they are White Dwarfs (de Ruiter et al., 2025; Hurley-Walker et al., 2024). While the remaining sample have no clear M-Dwarf companions, the recent discovery of a secondary $P=8.74$ hr period of GPM J1839-10 solely from radio timing suggests the White-Dwarf-M-Dwarf interpretation may extend to OIR-quiet candidates (Horváth et al., 2025b; Men et al., 2025). Nonetheless, a plethora of both neutron star and White Dwarf-M Dwarf binary models have been proposed to explain these new phenomena (e.g. Beniamini et al., 2023, 2020; Cooper and Wadiasingh, 2024; Qu and Zhang, 2025).
@@
Perhaps most promising is the discovery of four ASKAP LPRTs (ASKAP J1755-2527, ASKAP J1935+2148, ASKAP J1839-0756, ASKAP/DART-J1832-0911), LOFAR’s ILT J1101+5521, and the MWA’s GPM J1839-10 and GLEAM-X J0704-37 using second-duration radio image differencing pipelines (Wang et al., 2025b; Li et al., 2024; Lee et al., 2025; Caleb et al., 2024; de Ruiter et al., 2025; McSweeney et al., 2025; Dobie et al., 2024; Hurley-Walker et al., 2023). Radio imaging consists of binning the U-V (East-West and North-South coordinates of each antenna pair’s separation) visibility coordinates on a grid and applying a 2-dimensional FFT. This is much faster than visibility beamforming and is mathematically identical, though some resolution and flux sensitivity are lost via the gridding process. The image-plane LPRT detections demonstrate that this sensitivity loss is negligible for LPRT timescales (seconds-to-minutes), motivating the development of real-time LPRT image plane search pipelines like the Commensal Real-time ASKAP Fast-Transients Coherent (CRACO) Upgrade (Wang et al., 2025a), in addition to upcoming beamformer-based searches such as the CHIME/SLOW pipeline (J. Huang et al., in prep.).
In this paper, we introduce the Deep Synoptic Array’s (DSA-110) radio telescope’s Not-So-Fast Radio Burst (NSFRB) pipeline, an image-plane single-pulse search for LPRTs on 0.134-160.8 second timescales. The NSFRB system builds on the existing real-time Fast Radio Burst (FRB) search, which has discovered $\sim 60$ FRBs and localized $\sim 30$ to host galaxies with arcsecond-scale precision (Ravi et al., 2023; Law et al., 2024; Sharma et al., 2024). In Section II we describe the NSFRB system design and search pipeline; in Section III we describe the commissioning NSFRB injection tests and observations of pulsar B0329+54; in Section IV we report the results of a pilot DSA-110 NSFRB Galactic Plane Survey (DN-GPS); in Section V we discuss the sensitivity limits placed on LPRT emission in the Galactic Plane, and its implications for the White Dwarf binary model; in Section VI we conclude with next steps, specifically the real-time search which is currently underway.
@@ II NSFRB System Design
The DSA-110 is a 96-antenna drift-scan (see e.g. Mantovanini et al., 2025; Wayth et al., 2015, for examples of drift-scan transient surveys with the MWA) radio interferometer at the Owens Valley Radio Observatory (OVRO) in Big Pine, CA. Details on the 64-antenna deployment and commissioning FRB survey can be found in Ravi et al. (2023). The array was completed in early 2025 by supplementing the East-West row of 47 “core” antennas with a 35-antenna North-South arm, enabling real-time localization of radio sources without incorporating the 14 long-baseline “outrigger” antennas. The antenna design, RF signal chain, baseband processing, and general compute resource layout remain largely similar to the 64-antenna deployment, and will not be detailed here. The revised visibility correlator, bandpass calibration procedures, and the FRB search pipeline will be discussed in detail elsewhere (Ravi et al., in prep.) For this work, we focus on components specific to the NSFRB system. Code for the NSFRB search is public at https://github.com/dsa110/dsa110-nsfrb, all of which is implemented using Python, leveraging the JAX [^2] https://docs.jax.dev/en/latest/index.html library to deploy GPU tasks. An ETCD [^3] https://etcd.io/ server is used to manage alerts and coordination between compute nodes. Figure 1 shows the full real-time system diagram, which we describe in the next sub-sections.
@@ II.1 Fringe-Stopping and Imaging
for the DSA-110, $b_{\rm max}=2.6\,$km, $\lambda_{c}=20$ cm, where $\delta$ is the declination (e.g. Thompson et al., 2017). The array operates in drift-scan mode (no azimuthal slew) at $\delta=71.6^{\circ}$ during normal operations, making the nominal fringe rate $t_{F0}\approx 3.36\,$s.
Fringe-stopped visibilities are written to a 14.9 MB buffer read by the correlator node imager, which grids visibilities with baseline lengths $b>b_{\rm min}=20\,$m from the dense 82-antenna ‘core’ ($b<b_{\rm max}=485\,$m) to a $175\times 175$-pixel UV-grid. The core is assumed to be coplanar, so the W-term is neglected for normal operation; however, W-stacking is implemented and may be explored in the future (e.g. Gheller et al., 2023). Outrigger antenna data is included only for detailed manual inspection of candidates. The two visibility polarizations are summed, then approximately uniform weighting (Briggs robustness parameter $\mathcal{R}=-2$; Briggs, 1995) is applied to accumulate visibilities, and a 2-dimensional inverse-FFT [^4] The NumPy Python Package implementation of np.fft.ifft2d is used for imaging. The visibility grid is first shifted using np.fft.ifftshift to properly order the zero-frequency component; np.fft.ifftshift is then applied to the resulting image. is applied to form the final image with $\Delta\alpha=10.3^{\prime\prime}(\cos{\delta})^{-1}\,,\,\Delta\delta=10.3^{\prime\prime}$ pixel scale (3 pixels per $31^{\prime\prime}$ synthesized beamwidth). The DSA-110 primary beam FWHM is roughly $\Omega=1.5^{\circ}(\cos{\delta})^{-1}\times 1.5^{\circ}$, while the $175\times 175$ pixel image covers a $\Omega=0.5^{\circ}(\cos{\delta})^{-1}\times 0.5^{\circ}$ search region. Each of the 8 channels are imaged individually, then summed together to form a single image for each 134134\,ms time sample. Images on each node are then buffered (6.125 MB) for transport by an http client to a central search server.
@@
The custom cerberus single-pulse search pipeline implements Python JAX Just-In-Time (JIT) compiled functions to de-disperse and boxcar filter images on the GPU. Data is first median subtracted in each pixel, then de-dispersed over 16 DM trials between 1714000171-4000 pc cm-3 (DM tolerance of 1.6). The JAX de-dispersion method is adapted from a custom PyTorch implementation, PyTorchDedispersion [^5] https://github.com/nkosogor/PyTorchDedispersion (Paszke et al., 2019; Kosogorov et al., 2025), and allows shift indices for each frequency channel to be pre-computed and loaded onto the GPU to minimize latencies. The resulting de-dispersed time series are boxcar filtered on $\log_{2}$ spaced trials from 1-16 samples (0.1342.140.134-2.14 s), then normalized by the running standard deviation to compute the S/N. Final candidates are found by taking the maximum S/N (and its TOA) over time. A nominal $6\sigma$ signal-to-noise (S/N) threshold is imposed to identify candidates, which are pushed to the post-processor.
The search is performed at three sampling timescales: 0.134 s, 0.67 s (binned by 5 samples), and 10.05 s (binned by 75 samples). The process server bins images at each timescale as they are received, aligning them in the RA direction to account for the 25-sample fringe-stopping interval, and preserves batches for cerberus to re-search at the 0.67 s and 10.05 s timescales. For the real-time search, a single search iteration on the next shortest timescale (0.134 s, 0.67 s) is dropped to enable the longer timescale (0.67 s, 10.05 s) search iterations (the Galactic Plane data is not searched in real-time and thus does not drop iterations). The 0.67 s data is searched identically to the 0.134 s data, now probing DMs from 85520000855-20000\,pc cm-3666While it is not strictly necessary to search up to 2000020000\,pc cm-3 since the peak DMs in the Galactic Plane typically only reach $\sim 5000$pc cm-3, this range was selected because (1) the coarse time and frequency resolution, there is little benefit from finely sampled DM trials, and (2) extending the DM range by exactly $5\times$ for data sampled at $5\times$ lower resolution allows the GPU search code to utilize the same pre-computed shifts, thus saving GPU compute time and memory by only loading the the lookup table once. with widths from 0.6710.720.67-10.72\,s. The longest timescale data bypasses de-dispersion, leveraging the fact that the DM delay for DMs below $\sim 20000\,$pc cm-3 is now contained to a single sample across the 187.5 MHz band. They are then searched from 10.05160.810.05-160.8 s. The sensitivity of the search is shown in Figure 2, which indicates the full range of single-pulse widths the NSFRB search probes. The broad morphology of known LPRTs includes those with smooth profiles (e.g. McSweeney et al., 2025) and ‘spiky’ sub-pulses (e.g. Li et al., 2024; Dong et al., 2025b). By searching at multiple timescales, cerberus will be sensitive to both morphologies.
@@
Figure 2: Pulse width and $6\sigma_{\rm srch}$ Fluence Sensitivity to $\Delta t=134$ ms Bursts of the DSA-110 FRB (light blue; $\sim 840\,$mJy ms) and NSFRB (dark blue hatched; $\sim 40\,$Jy ms) searches. For each search we assume the fluence threshold is fixed; thus the lower border indicates the corresponding flux density threshold $6\sigma_{\rm th}\Delta t/W$ where $W$ is the width measured at 1.4 GHz (note that in some cases (e.g. ASKAP J1832-09 Li et al., 2024; Wang et al., 2025b), only sub-pulses were detected at 1.4 GHz while a broader profile was detected at lower frequencies. In these cases, we still use the 1.4 GHz sub-pulse width). The dark blue dashed line shows the 90% completeness limit $25\sigma\approx 1200$ mJy ($25\sigma\Delta t\approx 160$ Jy ms; see Section III.2). The purple dashed line shows the theoretical radiometer $6\sigma_{\rm th}$ sensitivity of the NSFRB survey. 1.4 GHz fluences and widths are shown for radio pulsars in the ATNF catalog (diamonds), FRBs from the literature (triangles), and LPRTs (circles). DSA-110 FRBs have dark blue outlines and show their detection pulse widths and S/N rather than refined estimates from baseband analysis. The colorbar indicates distance estimates in kpc (FRB luminosity distances are computed from host Galaxy redshifts; pulsar, magnetar, and LPRT distances are from parallaxes where available, and DM distance limits for the rest of the sample). Note that only pulsars, FRBs, and LPRTs with 1.4 GHz detections are shown.